# Learning Individual Behavior in Agent-Based Models with Graph Diffusion Networks

### v1

@@ 4 Experiments
In this section, we present different experiments to assess and demonstrate our framework’s ability to learn micro-level agent behaviors and faithfully reproduce emergent system-level dynamics. We evaluate our Graph Diffusion Network on two canonical agent-based models: the Schelling’s segregation model and the Predator–Prey ecosystem, presented in Section 2 and detailed in Supplementary Section B. We test both its micro-level and macro-level fidelity. At the micro level, we measure how well the surrogate reproduces the conditional next-state distribution of each agent under identical context on an out-of-training ramification data set. At the macro level, we assess whether the surrogate, once trained on the first $T_{\mathrm{train}}=10$ timesteps, can accurately reproduce the subsequent $T_{\mathrm{test}}=25$ timesteps of aggregate summary statistics.
Because no existing method directly accepts graph-structured agent states and outputs per-agent state distributions, we compare against an ablation variant in which the GNN embedding is replaced by a flat concatenation of all agent states (see Section 4.1). Both models are trained on the same ramified data and evaluated under identical protocols. In the remainder of this section, we first describe the experimental setup, including dataset generation, model variants, and evaluation metrics. We then present a qualitative analysis of emergent patterns, followed by a comprehensive quantitative comparison. All implementation and reproducibility details are provided in the Supplementary Materials. Full code to reproduce our experiments is available at the following link: https://github.com/fracozzi/ABM-Graph-Diffusion-Network/.
@@ 4.1 Experimental design
Our core hypothesis is that both relational structure and stochastic modeling are crucial for accurate ABM surrogates. We consider therefore two possible ablations. In the first, we remove the message-passing GNN and replace the interaction graph with a flat concatenation of all agents’ state vectors—this isolates the impact of neglecting agent interactions. The second would drop the diffusion component entirely, yielding a purely deterministic model incapable of capturing any randomness; however, since such a model cannot produce a stochastic output, it would offer little insights. Thus, we focus on the GNN-removed variant to justify modeling the interaction graph.
@@
We evaluate our approach on the two ABMs described in Section 2 as case studies. The first is the Schelling segregation model, in which $n$ agents occupy cells on a two-dimensional grid. Each agent has a fixed binary “color” and a position on the grid. At each timestep, an agent is considered happy if the proportion of its (up to eight) immediate neighbors sharing its color exceeds a tolerance threshold $\xi$; otherwise, it is unhappy and relocates to a randomly selected empty cell; thus, the interaction graph $G_{t}$ links each agent to its neighbors at time $t$. We adopt the standard NetLogo implementation of this model (wilensky1999netlogo, ). The second is a predator–prey ecosystem model, where agents belong to one of two species (predator or prey), inhabit grid cells, and cycle through life phases—Unborn, Alive, Pregnant, and Dead. At each timestep, an Alive agent may move to a neighboring cell, reproduce (becoming Pregnant), or die, with probabilities specified by a parameter matrix $\Psi$ and conditioned on the local presence of predators or prey (tang2022data, ; wilensky2015introduction, ). Pregnant agents revert to Alive after giving birth; Unborn agents become Alive if their parent is Pregnant; and Dead agents remain inactive. Here, $G_{t}$ links Alive neighboring agents, with specific rules for Pregnant and Unborn agents. See Supplementary Section B for more details. In both ABMs, each agent’s full state at time $t$ comprises its position, type (color or species), and, for the predator-prey ABM, its life phase. Together, these two models span both simple relocation dynamics and richer birth–death interactions, providing diverse testbeds for our surrogate.
@@
Figure 2: Evolution of the position of black and red agents in the Schelling model, for three simulations runs, one for each of the considered tolerance thresholds $\xi_{1}=0.625$, $\xi_{2}=0.75$, $\xi_{3}=0.875$. We compare the ground truth (top row) with our surrogate (middle row) and the ablation (bottom row), showing for each three time steps, $t=0$ (initial conditions, same for each column but kept for clarity), $t=15$, and $t=30$.
@@
To quantify how faithfully our surrogate captures individual agent behavior, we compare its predicted conditional next-state distributions against the ABM’s true stochastic transitions using the Earth Mover’s Distance (EMD) (EMD, ). We extend the ramification dataset beyond the training horizon and generate corresponding datasets for both the surrogate and the ablation models. The EMD is then computed as the mean value across timesteps and individual agents. In the Schelling ABM, the EMD compares the distribution of agent positions. This directly measures the surrogate’s ability to relocate unhappy agents correctly, while keeping happy agents stationary. In the predator-prey model we treat the agent’s categorical life phase as the random variable and compute the EMD over its four-state distribution. This metric captures both deterministic transitions (e.g., Unborn $\to$ Alive, Dead $\to$ Dead) and stochastic, interaction-driven transitions (e.g., Alive $\to$ Dead, Alive $\to$ Pregnant).
@@
Next, we test whether agent-level predictions translate into faithful reproduction of emergent, system-level behavior. For each model, we track a summary statistic over time: the number of happy agents in Schelling, and the number of active (i.e. Alive and Pregnant) agents in the predator–prey ecosystem. Reusing the same ramification branches as in training would offer little new information, since different stochastic branches from the same state tend to produce very similar macroscopic trajectories. Instead, we generate a fresh ensemble of main-branch simulations ($100$ independent runs) beyond the training horizon. We then compute the symmetric mean absolute percentage error (sMAPE) between the mean ground-truth trajectory and the mean surrogate-predicted trajectory across this ensemble, providing a quantitative measure of the surrogate’s ability to capture oscillations and steady-state behavior truly out-of-sample.
@@
We consider three parameter combinations $\xi$ for the Schelling ABM, each producing distinct segregation outcomes, and four $\Psi$ combinations for the predator-prey ABM, reflecting different oscillatory patterns in the population dynamics. For each ABM and parameter setting, we simulate $T_{\mathrm{train}}=10$ main-branch steps with $R=500$ stochastic branches per step, yielding the training ramification as in Figure 1. For macro-evaluation, we run $100$ independent main-branch simulations to calculate sMAPE. For micro-evaluation, we generate an out-of-sample ramification dataset of $T=25$ timesteps. We train both surrogate and ablation for $100$ epochs using Adam with learning rate $10^{-5}$ for the diffusion model, and Adam with learning rate $2\cdot 10^{-5}$ for the GNN, batch size equal to number of agents in the system, and diffusion hyper-parameters $\tau_{\max}=100$ (more information in Supplementary Section A).
@@
Figure 3: Forecasting macro-level summary statistics (here, the number of alive preys and predators over time), starting from the last condition seen in training, for 100 independent simulation runs, under configuration $\Psi_{1}$ (oscillations for both predators and preys, top) and $\Psi_{4}$ (oscillations only for predators, bottom). Left: original ABM simulations. Center: surrogate. Right: ablation. The dashed vertical line indicates the end of the training phase for the surrogate and ablation models.
@@ 4.2 Results
Let us first consider the Schelling ABM, under the configurations $\xi_{1}=0.625$, $\xi_{2}=0.75$, $\xi_{3}=0.875$. Figure 2 illustrates how the ground-truth ABM (top row) progresses from a randomized initialization to structured, segregated communities for the first two configurations $\xi_{1},\xi_{2}$, while it remains unsegregated for $\xi_{3}$. At the first two tolerance levels, in fact, the agents gradually self-organize into distinct clusters, with segregated communities clearly emerging around $t=20$ (see Supplementary Section C). The middle row represents the evolution of the system according to our surrogate model: we initialize the system with the same starting condition, and then we iterate giving the current state $\textbf{Z}_{t}$ to our model, and using one sample of the generated output as the next state $\textbf{Z}_{t+1}$. We observe that our surrogate exhibits a qualitatively similar pattern of cluster formation over time, distinct for each of the three configurations, while the ablation model (bottom row) fails to meaningfully relocate agents, largely maintaining a random configuration.
@@
Next, we consider Predator-Prey ecological model. Figure 3 overlays 100 trajectories of prey and predator populations starting from the same state at $T_{\text{train}}=10$, comparing the stochastic trajectories from the ground-truth model with the one obtained by our surrogate and by the ablation. For both configurations, the surrogate and the ablation model are trained only with the initial time steps (up to the dashed line in the plots). Under the parameter set $\Psi_{1}$, the ground-truth ABM (top-left plot) exhibits classical Lotka–Volterra oscillations: a rise in prey growth drives a delayed increase in predators, which then triggers prey decline and a subsequent predator decline. Under $\Psi_{4}$, instead, only predators show a rise and decay, while preys only decline (bottom-left plot). We observe that the surrogate (center column) accurately captures both the phase lag and amplitude of these oscillations, while the ablation (right) collapses to near-monotonic trends. We perform the same analysis for alternative parameterizations $\Psi_{2}$, $\Psi_{3}$ (included in Supplementary Section C) that show different types of dynamics, as the population of predators and/or preys may exhibit monotonic extinctions. In all cases, the surrogate faithfully reproduces monotonic declines or single-peak dynamics and the ablation fails. We also observe (figures in Supplementary Section  C) that the surrogate recreates the rich spatial patterns of predator–prey clusters, also seen in similar settings in evolutionary game theory nowak1992evolutionary .
@@
Figure 4: Errors obtained by the proposed approach (Surrogate) and by the naive baseline (Ablation) in four different tasks. In the first column, error is measured as the EMD between the true and predicted distribution of individual (micro-level) behavior, i.e. predicting the next state of each agent from the previous one. In the second column, error is measured as the difference (sMAPE) in system-level quantities, i.e. comparing the true values of the number of agents with a given state with the one predicted by our model when trained on a fraction of the initial time steps (as in Figure 3). In the first row, we test three configurations of the Schelling model; in the second row, we compare four configurations of the Predator-Prey model.
@@
For the Schelling model, we observe that, on the micro level, the surrogate’s mean EMD is lower than the ablation’s mean EMD in all cases. The differences between the surrogate and the ablation are less pronounced at the thresholds $\xi_{1}$ and $\xi_{3}$. At $\xi_{1}$ (few unhappy agents) behavior is almost entirely deterministic and agents rarely move, while at $\xi_{3}$ (almost all agents unhappy) behavior is uniformly random, so even a flat, “always-move” or “never-move” rule yields near-optimal predictions in these two cases. In contrast, at the intermediate threshold $\xi_{2}$, where roughly half the agents are unhappy, the difference between the surrogate’s and the ablation’s EMD is more pronounced. A similar pattern is observed in the macroscopic evaluation. The surrogate’s sMAPE remains below 0.2, whereas the ablation fails to distinguish happy from unhappy cases and suffers large macro-level errors. This gap confirms that only the full model, with explicit graph-based interaction modeling, can learn the conditional relocation rule critical in balanced regimes.
For the Predator–Prey model, regarding micro-level behavior, our surrogate achieves a low EMD from the ground truth on average, and it consistently outperforms the ablation model. These results confirm that our model is able to faithfully reproduce the complex dynamics of this ABM even at the individual agent-level (thus explaining the macro-level results shown by Figure 3). The most successful case is $\Psi_{2}$, where our surrogate exhibits a near-zero difference from the ground truth. This is explained by the fact that in this configuration, most agents follow deterministic update rules (e.g., dead $\rightarrow$ dead), which are perfectly recovered by our model—but not by the ablation. However, even when isolating only the stochastic transitions (e.g., alive $\rightarrow$ dead, alive $\rightarrow$ pregnant), the surrogate’s EMD remains below 0.09 across all parameter sets, reflecting its accurate capture of true randomness (see Supplementary Section C). By contrast, the ablation model’s EMD for these same stochastic rules averages around 0.2, often reaching 0.7–0.8. At the macro level as well, the surrogate consistently outperforms the ablation, generally achieving low error. The best result is obtained with $\Psi_{1}$, the most complex dynamic, where the surrogate achieves an average sMAPE of approximately 0.08. This configuration produces two distinct population peaks, and the surrogate faithfully reproduces both their timing and amplitude (Figure 3). The worst result is obtained with $\Psi_{2}$, as this configuration is almost monotonic and dominated by long, near-zero tails that are noisy at very small scales, making them difficult for any model to reproduce.
@@ 5 Discussion
Limitations. Our approach is limited by our assumptions on the characteristics of the ABM to emulate. First, the interaction graph is assumed to be fully known. Future works might remove this limitation by estimating such a graph directly from available data. However, the estimation of a latent interaction graph is a follow-up challenge, for which our GNN-based approach represents a necessary first step. Second, highly sophisticated ABMs may include features not addressed in our framework - such as all-to-all interactions, multiple rounds of decision-making, or sequential stochastic events within a single time step. Capturing these dynamics may require extending our architecture to incorporate sequential or hierarchical components. While our method may not yet fully generalize to such settings, our findings demonstrate that building surrogates capable of replicating individual-level behavior is both feasible and effective, laying the groundwork for broader applications.
Future work. Building on this foundation, the differentiability of our surrogate opens up a range of powerful applications. It enables the use of gradient-based methods for any optimization task such as policy optimization (agrawal2025robust, ). It allows for efficient calibration of macro parameters by treating key parameters as additional inputs to the neural network. Most importantly, our approach naturally allows for the estimation of micro (i.e., agent) level variables — a challenge for ABMs, that often requires the ad hoc development of handcrafted probabilistic models (monti2020learning, ; monti2023learning, ). In fact, our model already contains such parameters expressed as agents’ individual states ($\textbf{Z}^{(i)}_{t}$), something typically not available in ABM surrogates (glielmo2023reinforcement, ). In doing so, our framework helps make ABMs more data-driven and empirically grounded, with promising applications in several scientific domains, such as economics pangallo2024datadriven , epidemiology gozzi2025epydemix , sustainability lamperti2025complex , urban science birkin2025digital , and ecology stock2024plant .

### v2

@@ 4 Experiments
In this section, we present different experiments to assess and demonstrate our framework’s ability to learn micro‐level agent behaviors and faithfully reproduce emergent system‐level dynamics. We evaluate our Graph Diffusion Network on two canonical agent‐based models: the Schelling’s segregation model and the Predator–Prey ecosystem, presented in Section 2 and detailed in Supplementary Section B. We test both its micro‐level and macro‐level fidelity. At the micro level, we measure how well the surrogate reproduces the conditional next‐state distribution of each agent under identical context on an out-of-training ramification data set. At the macro level, we assess whether the surrogate, once trained on the first $T_{\mathrm{train}}=10$ timesteps, can accurately reproduce the subsequent $T_{\mathrm{test}}=25$ timesteps of aggregate summary statistics.
Because no existing method directly accepts graph‐structured agent states and outputs per‐agent state distributions, there are no directly applicable baselines for our approach. Existing surrogates typically operate at the macro level, learning mappings from parameters to aggregate outcomes rather than reproducing full system dynamics. However, such approaches, including standard time-series models such as AR(1), fail to capture non-monotonic or cyclic patterns (e.g., predator–prey oscillations) and do not generalize beyond in-sample dynamics, as we show in Supplementary Section C.5. Therefore, we evaluate against two ablated variants. The first replaces the GNN embedding with a flat concatenation of all agent states, removing relational structure. The second keeps the GNN but removes the diffusion component, predicting deterministic next states instead. Both ablations are trained on the same ramified datasets and under identical protocols.
In the remainder of this section, we first describe the experimental setup, including dataset generation, model variants, and evaluation metrics. We then present a qualitative analysis of emergent patterns, followed by a comprehensive quantitative comparison. All implementation and reproducibility details are provided in the Supplementary Materials. Full code to reproduce our experiments is available at http://github.com/fracozzi/ABM-Graph-Diffusion-Network.
@@ 4.1 Experimental design
Our core hypothesis is that both relational structure and stochastic modeling are crucial for accurate ABM surrogates. We consider therefore two possible ablations. In the first, we remove the message-passing GNN and replace the interaction graph with a flat concatenation of all agents’ state vectors—this isolates the impact of neglecting agent interactions. The second drops the diffusion component entirely, yielding a purely deterministic model similar in spirit to prior GNN-based approaches for deterministic automata grattarola2021learning. In this ablated version, we predict the next agent state by minimizing the mean squared error (MSE) with respect to the true next agent state. Details about the architecture can be found in Supplementary Section A. These ablations allow us to measure the improvement achieved by combining relational and stochastic modeling.
@@
We evaluate our approach on the two ABMs described in Section 2 as case studies. The first is the Schelling segregation model, in which $n$ agents occupy cells on a two-dimensional grid. Each agent has a fixed binary “color” and a position on the grid. At each timestep, an agent is considered happy if the proportion of its (up to eight) immediate neighbors sharing its color exceeds a tolerance threshold $\xi$; otherwise, it is unhappy and relocates to a randomly selected empty cell; thus, the interaction graph $G_{t}$ links each agent to its neighbors at time $t$. We adopt the standard NetLogo implementation of this model (wilensky1999netlogo). The second is a predator–prey ecosystem model, where agents belong to one of two species (predator or prey), inhabit grid cells, and cycle through life phases—Unborn, Alive, Pregnant, and Dead. At each timestep, an Alive agent may move to a neighboring cell, reproduce (becoming Pregnant), or die, with probabilities specified by a parameter matrix $\Psi$ and conditioned on the local presence of predators or prey (tang2022data; wilensky2015introduction). Pregnant agents revert to Alive after giving birth; Unborn agents become Alive if their parent is Pregnant; and Dead agents remain inactive. Here, $G_{t}$ links Alive neighboring agents, with specific rules for Pregnant and Unborn agents. See Supplementary Section B for more details. In both ABMs, each agent’s full state at time $t$ comprises its position, type (color or species), and, for the predator-prey ABM, its life phase. Together, these two models span both simple relocation dynamics and richer birth–death interactions, providing diverse testbeds for our surrogate.
@@
Figure 2: Evolution of the position of black and red agents in the Schelling model, for three simulation runs, one for each of the considered tolerance thresholds $\xi_{1}=0.625$, $\xi_{2}=0.75$, $\xi_{3}=0.875$ (left, center, and right panel). We compare the ground truth (top row) with our surrogate (middle row) and with the best-performing ablation (according to sMAPE, bottom row). For each panel and model, we show three time steps: $t=0$ (initial conditions, same for each column but kept for clarity), $t=15$, and $t=30$.
@@
To quantify how faithfully our surrogate captures individual agent behavior, we compare its predicted conditional next‐state distributions against the ABM’s true stochastic transitions using the Earth Mover’s Distance (EMD) (EMD). We extend the ramification dataset beyond the training horizon and generate corresponding datasets for both the surrogate and the ablation models. The EMD is then computed as the mean value across timesteps and individual agents. In the Schelling ABM, the EMD compares the distribution of agent positions. This directly measures the surrogate’s ability to relocate unhappy agents correctly, while keeping happy agents stationary. In the predator-prey model, we treat the agent’s categorical life phase as the random variable and compute the EMD over its four‐state distribution. This metric captures both deterministic transitions (e.g., Unborn $\to$ Alive, Dead $\to$ Dead) and stochastic, interaction‐driven transitions (e.g., Alive $\to$ Dead, Alive $\to$ Pregnant).
@@
Next, we test whether agent‐level predictions translate into faithful reproduction of emergent, system‐level behavior. For each model, we track a summary statistic over time: the number of happy agents in Schelling, and the number of active (i.e. Alive and Pregnant) agents in the predator–prey ecosystem. Reusing the same ramification branches as in training would offer little new information, since different stochastic branches from the same state tend to produce very similar macroscopic trajectories. Instead, we generate a fresh ensemble of main‐branch simulations ($100$ independent runs) beyond the training horizon. We then compute the symmetric mean absolute percentage error (sMAPE) between the mean ground‐truth trajectory and the mean surrogate‐predicted trajectory across this ensemble, providing a quantitative measure of the surrogate’s ability to capture oscillations and steady-state behavior truly out-of-sample.
@@
We consider three parameter combinations $\xi$ for the Schelling ABM, each producing distinct segregation outcomes, and four $\Psi$ combinations for the predator-prey ABM, reflecting different oscillatory patterns in the population dynamics. For each ABM and parameter setting, we simulate $T_{\mathrm{train}}=10$ main‐branch steps with $R=500$ stochastic branches per step, yielding the training ramification as in Figure 1. For macro‐evaluation, we run $100$ independent main‐branch simulations to calculate sMAPE. For micro-evaluation, we generate an out-of-sample ramification dataset of $T=25$ timesteps. We train both surrogate and ablations for $100$ epochs using Adam with learning rate $10^{-5}$ for the diffusion model and Adam with learning rate $2\cdot 10^{-5}$ for the GNN, batch size equal to number of agents, and diffusion hyper-parameters $\tau_{\max}=100$ (more information in Supplementary Section A).
@@
Figure 3: Forecasting macro-level summary statistics (here, the number of alive preys and predators over time), starting from the last condition seen in training, for 100 independent simulation runs, under configuration $\Psi_{1}$ (oscillations for both predators and preys, top) and $\Psi_{4}$ (oscillations only for predators, bottom). From left to right: original ABM simulations, surrogate, diffusion-only ablation, GNN-only ablation. The dashed vertical line indicates the end of the training phase for the surrogate and ablation models.
@@ 4.2 Results
Let us first consider the Schelling ABM, under the configurations $\xi_{1}=0.625$, $\xi_{2}=0.75$, $\xi_{3}=0.875$. Figure 2 illustrates how the ground-truth ABM (top row) progresses from a randomized initialization to structured, segregated communities for the first two configurations $\xi_{1},\xi_{2}$, while it remains unsegregated for $\xi_{3}$. At the first two tolerance levels, in fact, the agents gradually self‐organize into distinct clusters, with segregated communities clearly emerging around $t=20$ (see Supplementary Section C). The middle row represents the evolution of the system according to our surrogate model: we initialize the system with the same starting condition, and then we iterate giving the current state $\textbf{Z}_{t}$ to our model, and using one sample of the generated output as the next state $\textbf{Z}_{t+1}$. We observe that our surrogate exhibits a qualitatively similar pattern of cluster formation over time, distinct for each of the three configurations. Instead, the ablation models fail to meaningfully relocate agents. The best-performing one, according to sMAPE, the diffusion-only model, largely maintains a random configuration.
@@
Next, we consider the Predator-Prey ecological model. Figure 3 overlays 100 trajectories of prey and predator populations starting from the same state at the end of training, comparing the stochastic trajectories from the ground-truth model with those obtained by our surrogate and by the ablation. For both configurations, the surrogate and the ablation models are trained only with the initial time steps (up to the dashed line in the plots). Under the parameter set $\Psi_{1}$, the ground-truth ABM (top-left plot) exhibits classical Lotka–Volterra oscillations: a rise in prey growth drives a delayed increase in predators, which then triggers prey decline and a subsequent predator decline. Under $\Psi_{4}$, instead, only predators show a rise and decay, while preys only decline (bottom-left plot). We observe that the surrogate (second column) accurately captures both the phase lag and amplitude of these oscillations, while the diffusion-only ablation (third column) collapses to near-monotonic trends. The GNN-only ablation (fourth column), besides following a completely deterministic dynamic, completely diverges from the ground truth. The number of preys under $\Psi_{1}$, for instance, quickly reaches a plateau at a value of 750 (not shown), almost twice the real one. We perform the same analysis for alternative parameterizations $\Psi_{2}$, $\Psi_{3}$ (included in Supplementary Section C) that show different types of dynamics, as the population of predators and/or preys may exhibit monotonic extinctions. In all cases, the surrogate faithfully reproduces monotonic declines or single‐peak dynamics and both ablations fail. We also observe (figures in Supplementary Section  C) that the surrogate recreates the rich spatial patterns of predator–prey clusters, also seen in similar settings in evolutionary game theory nowak1992evolutionary.
@@
Figure 4: Errors obtained by the proposed approach (Surrogate) and by the naive baselines (Diffusion-only and GNN-only ablation models) in four different tasks. In the first column, error is measured as the EMD between the true and predicted distribution of individual (micro-level) behavior, i.e. predicting the next state of each agent from the previous one. In the second column, error is measured as the difference (sMAPE) in system-level quantities, i.e. comparing the true values of the number of agents with a given state with the one predicted by our model when trained on a fraction of the initial time steps (as in Figure 3). In the first row, we test three configurations of the Schelling model; in the second row, we compare four configurations of the Predator-Prey model.
@@
For the Schelling model, we observe that, on the micro level, the surrogate’s mean EMD is lower than the diffusion-only ablation’s mean EMD in all cases. The differences between the surrogate and the diffusion-only ablation are less pronounced at the thresholds $\xi_{1}$ and $\xi_{3}$. At $\xi_{1}$ (few unhappy agents) behavior is almost entirely deterministic and agents rarely move, while at $\xi_{3}$ (almost all agents unhappy) behavior is uniformly random, so even a flat, “always‐move” or “never‐move” rule yields near‐optimal predictions in these two cases. In contrast, at the intermediate threshold $\xi_{2}$, where roughly half the agents are unhappy, the difference between the surrogate’s and the diffusion-only ablation’s EMD is more pronounced. A similar pattern is observed in the macroscopic evaluation. The surrogate’s sMAPE remains below 0.2, whereas the diffusion-only ablation fails to distinguish happy from unhappy cases, resulting in large macro‐level errors. In the GNN-only ablated model, at the micro-level EMD increases proportionally with the level of stochasticity introduced by the parameters, with EMD values similar to the surrogate for $\xi_{1}$ (the most deterministic) and much higher for $\xi_{3}$ (the most stochastic). At the macro-level, we observe large errors, with sMAPE always higher than the surrogate. These gaps confirm that only the full model, with explicit graph‐based interaction modeling and stochasticity, can learn the conditional relocation rule critical in balanced regimes.
For the Predator–Prey model, regarding micro-level behavior, our surrogate achieves a low EMD from the ground truth on average, and it consistently outperforms both ablation models. These results confirm that our model is able to faithfully reproduce the complex dynamics of this ABM even at the individual agent-level (thus explaining Figure 3). The most successful case is $\Psi_{2}$, where our surrogate exhibits a near-zero difference from the ground truth. In fact, in this configuration, most agents follow deterministic update rules (e.g., dead $\rightarrow$ dead), which are perfectly recovered by our model, but not by the diffusion-only ablation — which also obtains worse results on stochastic rules, as shown in Supplementary Section C.4. Instead, the GNN-only ablation performs well only in those deterministic cases, but fails in all the others. At the macro level as well, the surrogate consistently outperforms the ablation, generally achieving low error. The best result is obtained with $\Psi_{1}$, the most complex dynamics, where the surrogate achieves an average sMAPE of approximately 0.08. This configuration produces two distinct population peaks, and the surrogate faithfully reproduces both their timing and amplitude (Figure 3). The worst result is obtained with $\Psi_{2}$, as this configuration is almost monotonic and dominated by long, near-zero tails that are noisy at very small scales, making them difficult for any model to reproduce.
@@ 5 Discussion
Limitations. Our approach is limited by our assumptions about the characteristics of the ABM to emulate. First, the interaction graph is assumed to be fully known. Future work might remove this limitation by estimating such a graph directly from available data. However, the estimation of a latent interaction graph is a follow-up challenge, for which our GNN-based approach represents a necessary first step. Second, highly sophisticated ABMs may include features not addressed in our framework - such as all-to-all interactions, multiple rounds of decision-making, or sequential stochastic events within a single time step. Capturing these dynamics may require extending our architecture to incorporate sequential or hierarchical components. While our method may not yet fully generalize to such settings, our findings demonstrate that building surrogates capable of replicating individual-level behavior is both feasible and effective, laying the groundwork for broader applications.
[…]